Mean Standard Deviation
Men 797 482
Women 660 414

Use these estimates as the mean m and standard deviation s for the U.S. populations for these age
groups. If we take a random sample of 40 men and 35 women, what is the probability of obtaining a
difference between sample means of 100 mg or more?
1
Expert's answer
2020-04-16T18:03:03-0400

"\mu_m=797\\
\mu_w=660\\
\sigma_m=482\\
\sigma_w=414\\
n_m=40\\
n_w=35"

We should find "P\{\overline{x}_m-\overline{x}_w\geq 100\}=1-P\{\overline{x}_m-\overline{x}_w< 100\}."


"\overline{x}=\overline{x}_m-\overline{x}_w\text{ is random variable}.\\
\mu=M(\overline{x}_m-\overline{x}_w)=M(\overline{x}_m)-M(\overline{x}_w)=\mu_m-\mu_w=\\
=797-660=137\text{ (here we used the theorem } \mu_{\overline{x}}=\mu)."

Random variables "\overline{x}_m" and "\overline{x}_w" are independent. So "D(\overline{x}_m-\overline{x}_w)=D(\overline{x}_m)+D(\overline{x}_w)=482^2+414^2=\\
=403720.\\
\sigma=\sigma(\overline{x}_m-\overline{x}_w)=\sqrt{403720}.\\
n=n_m+n_w=40+35=75.\\
\text{As } n\geq 30\text{ we can apply the consequence of Central Limit Theorem.}\\
\text{We will consider random variable } Z=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}.\\
Z\text{ will be asymptotically normal:}\\
\lim_{n\rightarrow\infty}P\{Z\leq z\}=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^ze^{-\frac{t^2}{2}}dt.\\
\text{So } P\{\overline{x}<100\}=P\{Z<\frac{100-\mu}{\sigma/\sqrt{n}}\}=P\{Z<\frac{100-137}{\sqrt{403720}/\sqrt{75}}\}=P\{Z<-0.5043\}\approx 0.3070.\\
P\{\overline{x}\geq 100\}=1-P\{\overline{x}<100\}\approx 1-0.3070=0.693."